Circular No. 7672

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

WZ SAGITTAE T. Kato, R. Ishioka, M. Uemura, K. Matsumoto, and H. Ohashi, Ky-

oto University; G. Masi, Ceccano, Italy; G. A. Good, Albuquerque, NM; J. Pietz, Erftstadt, Germany; and M. Moilanen, Nyrola Observatory, on behalf of the VSNET collaboration team, write: “Our time-resolved CCD photometry reveals that WZ Sge started to show prominent, highly structured oscillations (amplitude 0.45–0.50 mag) on July 23.76 UT. The oscillations have had a stable period of 0.056652 ± 0.000006 day between July 24.46 and 27.01, which is in striking agreement with the orbital period (0.05669 day) of WZ Sge. This finding confirms the development of ‘early superhumps’, as predicted by Ishioka et al. (IAUC 7669). For more details of the evolution of early superhumps, see http://www.kusastro.kyoto-

u.ac.jp/vsnet/DNe/wzsge01.html.”

H. Baba, National Astronomical Observatory of Japan; K. Sadakane, Osaka-Kyoiku University; and Y. Norimoto, Okayama Astrophysical Observatory (OAO), communicate: “We made time-resolved spectroscopic observations (range 400–500 nm) of WZ Sge between July 24.576 and 24.701 UT with the OAO 0.91-m telescope. Spectra taken at about the middle of the eclipses show that He II 468.6-nm is strongly in emission with a doublepeaked profile. The estimated peak separation is ∼ 1.4 nm, but the profile of the peaks varies with the orbital period. The C III-N III complex at 464.0 nm is also in double-peaked emission. These variations may indicate an actual evolutionary effect during the early stage of the outburst. On the other hand, Hβ shows both emission and absorption features like a P-Cyg profile, varying with the orbital period. Higher Balmer series Hγ and Hδ are seen as broad absorption features, as is He I 447.1-nm. The emissionline features resemble the beginning of the 1978 outburst (cf. IAUC 3311; Ortolani et al. 1980, A.Ap. 87, 31) taken four days after maximum, but they does not resemble those in the spectra taken by K. Ayani at Bisei during the rising stage (IAUC 7669), which shows no emission component around the wavelengths of He II and the C-N complex.”

Corrigendum. On IAUC 7669, line 18, for Oltorani et al. 1980, read Ortolani et al. 1980,

Visual magnitude estimates: July 24.923 UT, 8.7 (R. Bouma, Groningen, The Netherlands); 24.96, 8.0 (E. Broens, Mol, Belgium); 25.094, 8.6 (P. Schmeer, Bischmisheim, Germany); 25.573, 8.3 (A. Pearce, Nedlands, W. Australia); 25.956, 8.4 (H. Bengtsson, Hisigs Backa, Sweden); 25.986, 8.7 (Bouma); 26.885, 8.7 (A. Kammerer, Ettlingen, Germany).

2001 July 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green